OpenStudy (anonymous):

George makes large cookies. These cookies are 18 inches in diameter. How much icing will he need to cover the top of each cookie?

OpenStudy (anonymous):

area= pi* r^2 so a= pi * 9^2 or pi* 81 or 3.14*81 or 254.34 square inches